OpenStudy (anonymous):

PLEASE HELP Help Please being trying to find answer for 1 hour now :(. Determine the equation of the line, in point slope form, that runs through points (6,-5) and (-2,-3) A: y+5= - 1/4 x-6 B: y-6= - 1/4 (x+5) C y+5= - 1/4 (x-6)

OpenStudy (anonymous):

Duh.... solve it out. How stupid are you?

OpenStudy (anonymous):

Dang im sorry for askin and question everyone aint good at math please go somewhere else wit that please.

OpenStudy (kamibug):

Hey there! :) Welcome to Openstudy! ^.^ Point-Slope Form is: y-y1=m(x-x2) All your options have -1/4 for the slope so we don't need to worry about finding it. :P You can choose either of the two points to plug into the equation but it looks like here they want you to use (6, -5). -5 will be y1 and 6 will be x1. Just substitute it into the equation! :D Remember, since the 5 is negative, the minus sign turns into a +. Two negatives make a positive! But, the 6 is positive, so we keep the minus sign. :) y+5 = -1/4(x-6)

OpenStudy (anonymous):

Thank you so much god bless you :)

OpenStudy (kamibug):

No problem! :)